find the measure of an angle if its measure is 50 degrees more than its complement.
Answer by DrBeeee(684)

Let x = one angle
Let y = x's complement
Then we have
(1) x + y = 90
We are given that x is y +50 (or y = x+50, it doesn't matter) so we get
(2) y + 50 + y = 90 or
(3) 2*y = 40 or
(4) y = 20
Then
(5) x = y + 50 or
(6) x = 20 + 50 or
(7) x = 70
Let's check this with (1).
Is (70 + 20 = 90)?
Is (90 = 90)? Yes
Answer: The angle is 70 degrees.